OpenStudy (anonymous):

Can anyone help me out. f(x)=1over 8x+9. Thanks!

OpenStudy (anonymous):

What exactly are you suppose to do with the function?

OpenStudy (anonymous):

Find the inverse of the function. Here are the possible solutions: A) f-1(x)=8x+72 B) f-1(x)=1over8x-9 C) f-1(x)=1 over 9x+8 D) f-1(x)=8x-72

OpenStudy (anonymous):

I'm stumped. Is this what you're trying to do? http://www.tutorvista.com/math/inverse-of-polynomial-function

OpenStudy (anonymous):

Is it \[f(x) = [1/8x]+9\] OR \[f(x) = 1/(8x+9)\] There's a difference when solving both.

OpenStudy (anonymous):

PS. The first of Veronometry's equations will give you a match. Follow the steps in the link above.

OpenStudy (anonymous):

thank your. I will

OpenStudy (anonymous):

I would be X is an element of (-infinite, -9/8) or (-9/8, infinite) Because you set the bottom, Underneath the bar, to equal zero... Because division of 0 is undefined. Then what ever your answer is is the one thing that X cannot be.